UN_5 Review questions 

1. Discuss the main assumptions of perfectly competitive market

2. Describe the feature of monopolistic competition that resembles perfect competitive and 

the monopolistic market structure.

3. What is the difference between real and fancied differentiation. Explain using practical

examples.

4. What are the similarities and differences between oligopoly and monopolistically 

competitive market structure?

5. A firm operates in a perfectly competitive market. The market price of its product is 4 

birr and the total cost function is given by TC= 1/3Q^3-5Q^2+20Q+50 ,where TC is the 

total cost and Q is the level of output.

a) What level of output should the firm produce to maximize its profit? 

b) Determine the level of profit at equilibrium. 

c) What minimum price is required by the firm to stay in the market?



1
Expert's answer
2021-04-19T07:49:32-0400

1. The basic assumptions are: the product is homogeneous (same or identical products), there are many buyers and sellers, consumers have perfect information, and there are no barriers to entry or exit (easy entry and exit).

2. Monopolistic competition has a lot of buyers like in perfect competition, but the sellers have some amount of market power like in a monopoly.

3. It is real when there are slight differences in the product of the firm as in taste if it is a foodstuff, or in quality etc. It is fancied if the difference is just to attract the customer and not real differentiation, for example, differences in packaging, design etc.

4. Oligopoly is a market structure containing a small number of relatively large firms, with significant barriers to entry of other firms. Monopolistic competition is a market structure containing a large number of relatively small firms, with relative freedom of entry and exit.

The similarities between oligopoly and monopoly competition are: They both exhibit imperfect competition in that oligopoly has few sellers while monopoly has many sellers. Firms have some level of control over prices in both competitive structures.

5. a) The level of output the firm should produce to maximize its profit is when MC = P,

"Q^2 - 10Q + 20 = 4,"

(Q - 8)(Q - 2) = 0,

Q1 = 8 units, Q = 2 units.

b) The level of profit at equilibrium (Q = 8) is:

"TP = 4\u00d78 - (1\/3\u00d78^3-5\u00d78^2+20\u00d78+50) = -28.67."

c) Minimum price required by the firm to stay in the market is when P = AVC.